# DevBench: A Comprehensive Benchmark for Software Development

### 13 Mar 2024 v1

@@ 1 Introduction
Given its practical value and reasoning challenges, programming has become an important domain to deploy and evaluate large language models (LLMs), leading to popular products like GitHub Copilot and research benchmarks like HumanEval (Chen et al., 2021) and APPS (Hendrycks et al., 2021). While these earlier coding tasks focused on generating a single code file or even a single method conditioned on simple instructions, recent works such as SWE-bench (Jimenez et al., 2023) and RepoBench (Liu et al., 2023b) propose new evaluations that challenge LLMs with repository-level coding tasks, which feature longer, more involved NL2Code problems. that capture practical challenges in the real world. Still, these benchmarks concentrate on narrow aspects of software development. To this end, datasets and metrics that comprehensively evaluate software development as a multi-phase set of tasks are notably absent.
@@
We also develop a baseline system by employing ChatDev (Qian et al., 2023a, b) as the foundational framework to benchmark state-of-the-art pretrained code models. We find that all models fail to solve the challenges presented within DevBench. GPT-4-Turbo achieves the highest scores amongst all evaluated models, yet it obtains less than 10% on our repository-level implementation task. Other tasks prove relatively more manageable for models; however, even GPT-4-Turbo struggles to attain scores exceeding 40% on any given task. For instance, models generally fail to generate executable testing code, with oracle test scores falling below 40%. Despite this, the generated testing code demonstrates potential in terms of coverage score, achieving as high as 79.4% when it is executable. Furthermore, our investigation into different prompting methods shows that only prompts supplemented with external information, execution feedback in our context, yield notable and consistent improvements. We employ LLM-as-a-Judge (Zheng et al., 2023a) for evaluating software design, where the outcomes exhibit high agreement with human annotations. Importantly, qualitative analysis shows that models demonstrate difficulties in handling Makefile and Gradle, configuring function arguments and employing advanced object-oriented programming techniques. Overall, our DevBench introduces a novel challenge for existing code models, and our investigation sheds light on fundamental issues, paving the way for future research.
@@ 3.1 Task 1: Software Design
During this phase, the model is tasked with interpreting the Product Requirements Document (PRD) to create Unified Modeling Language (UML) diagrams, utilizing the extensively recognized Mermaid syntax, alongside formulating architectural designs using hierarchical file-tree structures. The class diagram acts as a fundamental component in software system modeling, illustrating classes, their attributes, and interrelations. Sequence diagrams detail processes and objects involved and the sequence of messages exchanged to carry out the software functionality. Architectural design is aimed at crafting a structured layout of source code, compilation scripts, and auxiliary files crucial for software implementation. For exemplifications of UML diagrams and architectural frameworks, refer to Appendix A.1.
Evaluation. Since the Software Design tasks are open-ended, we employ the LLM-as-a-judge approach (Zheng et al., 2023a; Wang et al., 2023; Chiang & Lee, 2023) to conduct the automatic evaluation. The evaluation is anchored by two principal metrics: general principles and faithfulness. Detailed evaluation guidance is found in Appendix LABEL:subsubsec:preference_metrics.
The general principles metric plays a crucial role, with each task sharing common elements while maintaining specific criteria. For all three sub-tasks, principles like cohesion and decoupling, and practicability are fundamental. Cohesion and decoupling emphasize the importance of clarity and functionality within individual elements (classes or sequences) and reducing dependencies between different components. In terms of practicability, all tasks require designs to be readable, understandable, and modular, facilitating ease in development, testing, and maintenance. Meanwhile, each task has its unique focus areas: UML Class diagrams are evaluated on complexity; UML Sequence diagrams concentrate on uniformity, integration, and interaction complexity; Architecture Design highlights the distinction between design and coding, and conformance to practical standards. Subsequently, the faithfulness metric gauges the extent to which models adhere to specified instructions.
@@ 4.1 Dataset Construction
The final phase, document preparation, involved the creation of standard software design documents, namely, UML Class and Sequence Diagrams, and Architecture Designs, for each repository. We provide annotators with specific guidelines and templates for these documents. The annotators were responsible for ensuring that these design documents corresponded accurately and cohesively with the respective code repositories.

### 15 Mar 2024 v2

@@ 1 Introduction
Given its practical value and reasoning challenges, programming has become an important domain to deploy and evaluate large language models (LLMs), leading to popular products like GitHub Copilot and research benchmarks like HumanEval (Chen et al., 2021) and APPS (Hendrycks et al., 2021). While these earlier coding tasks focused on generating a single code file or even a single method conditioned on simple instructions, recent works such as SWE-bench (Jimenez et al., 2023) and RepoBench (Liu et al., 2023b) propose new evaluations that challenge LLMs with repository-level coding tasks, which feature longer, more involved NL2Code problems, that capture practical challenges in the real world. Still, these benchmarks concentrate on narrow aspects of software development. To this end, datasets and metrics that comprehensively evaluate software development as a multi-phase set of tasks are notably absent.
@@
We also develop a baseline system by employing ChatDev (Qian et al., 2023a, b) as the foundational framework to benchmark state-of-the-art pretrained code models. We find that all models fail to solve the challenges presented within DevBench. GPT-4-Turbo achieves the highest scores amongst all evaluated models, yet it obtains less than 10% on our repository-level implementation task. Other tasks prove relatively more manageable for models; however, even GPT-4-Turbo struggles to attain scores exceeding 40% on any given task. For instance, models generally fail to generate executable testing code, with oracle test scores falling below 40%. Despite this, the generated testing code demonstrates potential in terms of coverage score, achieving as high as 79.4% when it is executable. Furthermore, our investigation into different prompting methods shows that only prompts supplemented with external information, execution feedback in our context, yield notable and consistent improvements. We employ LLM-as-a-Judge (Zheng et al., 2023a) for evaluating software design, where the outcomes exhibit high agreement with human annotations. Importantly, qualitative analysis shows that models demonstrate difficulties in handling Makefile and Gradle, configuring function arguments and employing advanced object-oriented programming techniques. Overall, DevBench introduces a novel challenge for existing code models, and our investigation sheds light on fundamental issues, paving the way for future research.
@@ 3.1 Task 1: Software Design
During this phase, the model is tasked with interpreting the Product Requirements Document (PRD) to create Unified Modeling Language (UML) diagrams, utilizing the extensively recognized Mermaid syntax, alongside formulating architecture designs using hierarchical file-tree structures. The class diagram acts as a fundamental component in software system modeling, illustrating classes, their attributes, and interrelations. Sequence diagrams detail processes and objects involved and the sequence of messages exchanged to carry out the software functionality. Architecture design is aimed at crafting a structured layout of source code, compilation scripts, and auxiliary files crucial for software implementation. For exemplifications of UML diagrams and architectural frameworks, refer to Appendix A.1.
Evaluation. Since the Software Design tasks are open-ended, we employ the LLM-as-a-judge approach (Zheng et al., 2023a; Wang et al., 2023; Chiang & Lee, 2023) to conduct the automatic evaluation. The evaluation is anchored by two principal metrics: general principles and faithfulness. Detailed evaluation guidance is found in Appendix B.
The general principles metric plays a crucial role, with each task sharing common elements while maintaining specific criteria. For all the sub-tasks, principles like cohesion and decoupling, and practicability are fundamental. Cohesion and decoupling emphasize the importance of clarity and functionality within individual elements (classes or sequences) and reducing dependencies between different components. In terms of practicability, all tasks require designs to be readable, understandable, and modular, facilitating ease in development, testing, and maintenance. Meanwhile, each task has its unique focus areas: UML Class diagrams are evaluated on complexity; UML Sequence diagrams concentrate on uniformity, integration, and interaction complexity; Architecture Design highlights the distinction between design and coding, and conformance to practical standards. Subsequently, the faithfulness metric gauges the extent to which models adhere to specified instructions.
@@ 4.1 Dataset Construction
The final phase, document preparation, involved the creation of standard software design documents, namely, UML Class and Sequence Diagrams, and Architecture Designs, for each repository. We provide annotators with specific guidelines and templates for these documents. The annotators were responsible for ensuring that these design documents corresponds accurately and cohesively with the respective code repositories.